OpenStudy (anonymous):

cos 2x + 3cossx + 2 =0 find th exact solution of the given equation in the interval [0,2pi)

OpenStudy (campbell_st):

well I'll assume its \[\cos^2(x) + 3\cos(x) + 2 = 0\] this can be factored to \[(\cos(x) +1)(\cos(x) + 2) = 0\] now some each for the factors for x.